OpenStudy (anonymous):

Find the length of the radius of a sphere with a surface area of 78.54 km2. A. 5 ft B. 15 ft C. 10 ft D. 2.5 ft

OpenStudy (turingtest):

what google didn't work this time?

OpenStudy (jamesj):

\[ Surface \ Area = 4 \pi r^2 \] where r is the radius. Use this formula to find the answer.

OpenStudy (anonymous):

Google is not smart enough.

OpenStudy (turingtest):

\[A=4\pi r^2\]so\[r=\sqrt{\frac{A}{4\pi}}\]so plug in the value you are given for the area A, out will come r, your answer, as James will testify

OpenStudy (anonymous):

k, thanks

OpenStudy (turingtest):

what did you get?

OpenStudy (anonymous):

well calculating things is actually hard; no ?

OpenStudy (turingtest):

It would seem

OpenStudy (anonymous):

wait, what the heck is the radius?

OpenStudy (anonymous):

The answer.

OpenStudy (anonymous):

yeah it's a heck! what the heck is that!!!

OpenStudy (anonymous):

That is what needs to be figured out. So, call it r for now.

OpenStudy (earthcitizen):

Just explain the damn thing!

OpenStudy (anonymous):

oh wait, sorry. that was a dumb question. I was looking at it backwards

OpenStudy (anonymous):

What's wrong with Google?

OpenStudy (turingtest):

your mantra:\[A=4\pi r^2\]so\[r=\sqrt{\frac{A}{4\pi}}\]

OpenStudy (anonymous):

EarthCitizen - Lot of people dedicate their time to help here. I don't think making demands on them and throwing arrogant insults is the best way to show respect and gratitude for their contributions.

OpenStudy (anonymous):

lol Mantra.

OpenStudy (earthcitizen):

whose arrongant and pompous by teasing ppl for the unfamiliarity with a topic! Focus!

OpenStudy (turingtest):

where is saifoo? he is neglecting his modship

OpenStudy (anonymous):

I don't remember anyone making fun of people for not knowing. Nobody even makes fun of not getting it. Yes, sometimes people do expect "learning" and insist on it.

OpenStudy (anonymous):

wait, you guys.. im totally lost :-/ i need help.

OpenStudy (anonymous):

pro-tempo-re mods can't be assigned much of tasks

OpenStudy (turingtest):

Just for good measure\[A=4\pi r^2\]so\[r=\sqrt{\frac{A}{4\pi}}\]plug in the surface area (given in the problem) where the A in the formula goes. Out comes r, the answer.

OpenStudy (earthcitizen):

then tech the best way YOU CAN!

OpenStudy (anonymous):

On the contrary I find EarthCitizen a bit swanky.

OpenStudy (earthcitizen):

teach

OpenStudy (anonymous):

TuringTest: May be she needs helps in finding the square root too ?

OpenStudy (anonymous):

Surface are of sphere = 4*pi*r^2 = 78.54km^2 So, r^2 = 78.54/(4*pi) So, r = SQRT(78.54/(4*pi)) So, r = ?

OpenStudy (anonymous):

2.5?

OpenStudy (turingtest):

Yes Yes Yes

OpenStudy (earthcitizen):

yep

OpenStudy (anonymous):

Yay! Good job.

OpenStudy (anonymous):

Congratz.

OpenStudy (anonymous):

oh thank god!

OpenStudy (anonymous):

Magnum opus!

OpenStudy (anonymous):

sorry you guys for being so cranky. along with math, I am also completing an essay. lots of work and multi tasking

OpenStudy (anonymous):

Apology accepted.

OpenStudy (anonymous):

We are thick skinned! Glad to see you do it.

OpenStudy (bimbim2697):

GT CAN U HELP ME PLZ?

OpenStudy (anonymous):

thankyou. still a lot more!! have the patience?

OpenStudy (anonymous):

Well, I am also dark skinned on top!! hahahahaahaha.

OpenStudy (anonymous):

lol! I ain't dark skinned like you GT :P

OpenStudy (earthcitizen):

i see the light in U!

OpenStudy (anonymous):

First of all question is somewhat wrong!!!!! Radius is a quantity representing length and radius doesn't need to be written as length of the radius. And coming to answer Surface area of sphere is 4*pi*r*r. From that r can be found out substituting others.